# Discovering Knowledge-Critical Subnetworks in Pretrained Language Models

### v1

@@ Abstract
Pretrained language models (LMs) encode implicit representations of knowledge in their parameters. However, localizing these representations and disentangling them from each other remains an open problem. In this work, we investigate whether pretrained language models contain various knowledge-critical subnetworks: particular sparse computational subgraphs responsible for encoding specific knowledge the model has memorized. We propose a multi-objective differentiable weight masking scheme to discover these subnetworks and show that we can use them to precisely remove specific knowledge from models while minimizing adverse effects on the behavior of the original language model. We demonstrate our method on multiple GPT2 variants, uncovering highly sparse subnetworks (98%+) that are solely responsible for specific collections of relational knowledge. When these subnetworks are removed, the remaining network maintains most of its initial capacity (modeling language and other memorized relational knowledge) but struggles to express the removed knowledge, and suffers performance drops on examples needing this removed knowledge on downstream tasks after finetuning.
@@ 1 Introduction
Large-scale pretrained language models (LLMs) encode large amounts of relational knowledge (Petroni et al., 2019; Carlini et al., 2023; Liu et al., 2023), which they adapt to successfully transfer to downstream tasks (Wang et al., 2019b; a). Due to this success, considerable prior research has focused on better understanding the extent to which LLMs capture different types of knowledge that are necessary for these tasks (Liu et al., 2019; Safavi & Koutra, 2021; Da et al., 2021; Huang et al., 2022). In these works, models are prompted using natural language verbalizations of relational triplets, which associate head and tail entities. Tokens in the sequence representing entities (or relations) in the triplets are masked, and the model must infill or complete the sequence to demonstrate it encodes the knowledge expressed by the sequence (Bosselut et al., 2019; Jiang et al., 2020).
Despite the body of work in studying LLMs as knowledge bases, less work has focused on where and how this knowledge may be encoded by the models that capture it. The answer to these questions could potentially facilitate the development of more effective finetuning methods, which can be useful for rectifying factual errors made by language models, keeping models up to date with evolving knowledge, and preventing ethically undesirable behavior. Works in probing (Belinkov & Glass, 2019; Durrani et al., 2020; Antverg et al., 2022; Belinkov, 2022) and mechanistic interpretability (Geva et al., 2021; 2022b; 2022a) discover hidden representations, neurons, and layers that are responsible for the expression of knowledge from these systems, but typically do not localize the knowledge accessing behavior at the weight-level. Recent work in model editing investigates whether specific knowledge in the model can be changed (De Cao et al., 2021; Dai et al., 2022; Hase et al., 2023b; Mitchell et al., 2022a; b; Meng et al., 2022; 2023; Hase et al., 2023a; Gupta et al., 2023; Jang et al., 2023; Chen et al., 2023). However, the goal of these methods is typically not to precisely localize the parameters responsible for encoding the knowledge, but instead to coarsely edit model parameters such that a new desired behavior (or knowledge) overwrites the model’s preference for the old one.
@@
Figure 1: We hypothesize the existence of knowledge-critical subnetworks that are responsible for expressing target knowledge triplets (TargetKG). When knowledge-critical subnetworks are removed, the remaining model can no longer express the specific triplets, but maintains its ability to express other relational knowledge (ControlKG) and its language modeling abilities (ControlLM). The lighter shades of blue illustrate neurons that lose weight connections in this process.
In this work, we hypothesize that language models contain particular sparse computational subnetworks that are responsible for expressing specific knowledge relationships. We call these subnetworks knowledge-critical as they are necessary for the model’s ability to express particular relational knowledge. As a result, when the knowledge-critical subnetwork is removed, the model’s ability to express the knowledge it represents is also removed, as represented by the remaining blue model in Figure 1 that can no longer correctly predict “restaurant” as the continuation of “A cafe is a type of ___”.
@@
To discover knowledge-critical subnetworks, we propose a multi-objective differentiable weight masking method over the original pretrained model. The remaining unmasked model loses the ability to express the target knowledge on which the mask was trained, but maintains its performance on other behaviors, thereby identifying the knowledge-critical subnetwork as the masked portion of the original model. We combine multiple objectives designed to (1) suppress the expression of target knowledge triplets, (2) maintain an ability to express generic relational knowledge, (3) maintain standard language modeling performance, and (4) encourage the subnetwork to be as sparse as possible. Combined, these objectives optimize a mask that promotes the removal of target knowledge, while maintaining the other behaviors of the pretrained language model.
@@
Our results — across multiple target knowledge graphs (constructed from WordNet and ConceptNet) and LLMs at multiple scales (from the family of GPT2 models) — show that our masking method consistently identifies sparse subnetworks ($\approx$98.6average parameters pruned) that satisfy our objectives. When these subnetworks are removed, the remaining model’s perplexity on the target knowledge associated with the subnetwork largely increases (on average, a relative perplexity increase of 257% for GPT2-small, 253% for GPT2-medium, and 5589% for GPT2-large), indicating that the expression of the target knowledge is successfully suppressed. The remaining network’s ability to model generic relational knowledge and natural language negligibly changes compared to the original model, implying the model maintains its original abilities. Finally, in a study on CommonsenseQA, we demonstrate that once these subnetworks are removed, models finetuned using parameter-efficient methods struggle with questions that require the knowledge encoded by the subnetwork.
@@ 2 Related Work
LLMs as Knowledge Bases   Our work builds on prior research that demonstrates the memorization of large-scale language models (LLMs) pretrained on massive amounts of web data (Carlini et al., 2021; AlKhamissi et al., 2022; Carlini et al., 2023). Multiple studies have depicted the different types of knowledge encoded by LLMs, including linguistic (Liu et al., 2019; Chen & Gao, 2022), relational (Safavi & Koutra, 2021), commonsense (Da et al., 2021), and actionable knowledge (Huang et al., 2022). Parametric knowledge in LMs is typically accessed in two ways. In the first, the model is conditioned with a natural language context and must complete or infill the sequence to identify the knowledge (Petroni et al., 2019; Liu et al., 2023; Yu et al., 2023). In these studies, human-defined discrete prompts and automatic prompt engineering are used to extract single and multi-token answers from language models (Jiang et al., 2020; Shin et al., 2020; Cao et al., 2021a; Zhong et al., 2021; Qin & Eisner, 2021). Alternatively, other methods fine-tune parameters to create an interface for accessing parametric knowledge (Bosselut et al., 2019; Roberts et al., 2020; Jiang et al., 2021; Hwang et al., 2021). In contrast, our work dives deeper into where knowledge is encoded by LLMs and proposes an algorithm to discover the subnetworks responsible for expressing these facts.
@@
Function-Specific Subnetworks   Methodologically, our work draws inspiration from work that identifies task-specific subnetworks in neural networks. Perhaps most known, Frankle & Carbin (2019) proposed the Lottery Ticket Hypothesis, which showed that learned subnetworks could achieve test accuracy similar to that of original networks. Other works pruned subnetworks for the purpose of efficient finetuning (Mallya et al., 2018; Zhao et al., 2020; Sanh et al., 2020; Guo et al., 2021), or identifying function-specific subnetworks (Cao et al., 2021b; Sanh et al., 2020; Zhang et al., 2021; Csordás et al., 2021). Identifying function-specific subnetworks also leads to useful applications, such as disentangling representations to reduce model susceptibility to spurious correlations (Zhang et al., 2021), probing models for linguistic properties (Cao et al., 2021b; De Cao et al., 2020), and finding subnetworks specialized for different languages (Foroutan et al., 2022). Most similar to our work is that of Ren & Zhu (2022), which learned coarse subnetworks that encoded large portions of ConceptNet. Similarly to these methods, we adopt a differentiable weight masking scheme, but use it to identify highly sparse subnetworks responsible for particular expressions of knowledge.
@@
Mechanistic Interpretability of LLMs   Mechanistic interpretability tackles the problem of understanding model behavior by reverse-engineering computations performed by transformer models. Elhage et al. (2021) discovered algorithmic patterns and frameworks in simplified transformer models. Following this framework, researchers discovered induction heads (Olsson et al., 2022), i.e., specific attention heads that can be the mechanistic source of general in-context learning in LLMs. Similarly, with interventions on multi-head self-attentions and MLP sublayers, Geva et al. (2023) identified two critical points where the model propagates information for predictions and the internal mechanism for attribute extraction. Another line of work focuses on knowledge tracing and localization in model parameters for the goal of model editing (Dai et al., 2022; Meng et al., 2022; 2023; Gupta et al., 2023; Hernandez et al., 2023). Activation patching with corrupted tokens (Meng et al., 2022) or corrupted prompts (Wang et al., 2023) use causal intervention to identify model activations responsible for flipping the model’s output. In contrast, our work focuses on preserving the original model to precisely locate model weights responsible for expressing a given set of target knowledge without counterfactuals. Our work is closer to path patching (Goldowsky-Dill et al., 2023) and automatic circuit discovery (Conmy et al., 2023), which focus on localizing behaviors to network subgraphs, but focuses specifically on identifying subnetworks associated with knowledge relationships.
3 Background
@@
To find a knowledge-critical subnetwork in a pretrained language model, we learn a differentiable weight mask (§4) over the parameters of the LM using a knowledge prediction task where a language model is prompted for relational knowledge.
@@
Prompting Language Models with Knowledge Graphs   We define a global relational KG as the set of knowledge triplets $K=\{(h_{1},r_{1},t_{1}),...,(h_{k},r_{k},t_{k}),...,(h_{n},r_{n},t_{n})\}$ where $h$ and $t$ are head and tail entity nodes, respectively, and $r$ is the relation that holds between the two entities. To input relational knowledge into a language model, triplets must be verbalized by instantiating a natural language template with the triplet components. For example, the knowledge triplet (house, IsA, building), can be reformulated with the IsA relation-specific template ‘‘{article} {$h$} is {article} {$t$}’’ as ‘‘A house is a building’’ A typical way to prompt for knowledge is to mask the tail entity ‘‘A house is a ___’’ (Petroni et al., 2019). Thus, to approximate an autoregressive model’s confidence on a given triplet, we can compute a distribution over the missing token and calculate the perplexity of the actual correct token building.
@@
Differentiable Weight Masking for Function-Specific Parameter Search   To localize parameters that are critical for modeling specific knowledge, we learn a binary mask over each network parameter. Consider a language model $f(x,{\bm{\theta}}){}$ with pretrained parameters ${\bm{\theta}}$ that takes as input $x$. We learn a set of binary parameters ${\bm{m}}\in\{0,1\}^{|{\bm{\theta}}|}$ that is element-wise multiplied with the frozen ${\bm{\theta}}$, such that our network is reformulated as $f(x,{\bm{m}}\odot{\bm{\theta}}){}$. Similar to other binary mask learning methods (Cao et al., 2021b; Sanh et al., 2020; Zhang et al., 2021), our method models each parameter mask ${\bm{m}}_{i}$ with the hard-concrete or gumbel-softmax distribution, a differentiable approach to learning continuous mask scores ${\bm{s}}_{i}\in[0,1]$ from real-valued parameters ${\bm{l}}_{i}\in{\mathbb{R}}$ (Maddison et al., 2017; Jang et al., 2017):
@@
where ${\mathcal{U}}_{1},{\mathcal{U}}_{2}\sim{\mathcal{U}}(0,1)$ and $\sigma$ is a sigmoid function. We use the approach of Csordás et al. (2021), which uses a straight-through estimator that thresholds the continuous score (Bengio et al., 2013):
@@ 4 Methodology
This section defines our methodology for finding knowledge-critical subnetworks with differentiable weight masking. We define our criteria for such a subnetwork and propose objectives that can optimize for the criteria.
Notation   We define a subnetwork as in §3: $f(x,{\bm{m}}\odot{\bm{\theta}})$, where ${\bm{\theta}}$ is the set of parameters of the network $f$ and ${\bm{m}}$ is the mask over a portion of that network’s parameters. We assume a target set of knowledge $K_{T}\subset K$ (TargetKG) for which we want to identify the responsible parameters.
@@ 4.1 Knowledge-Critical Subnetworks
Our overall goal is to find knowledge-critical subnetworks, which are essential parameters to express a given set of target knowledge. When knowledge-critical subnetworks are removed, the expression of the target triplets should be suppressed, and the expression of irrelevant triplets should be unaffected.
Suppression   For $f(x,{\bm{m}}\odot{\bm{\theta}})$ to be critical in expressing $K_{T}$, its removal from the original network should also remove the model’s ability to express the knowledge in $K_{T}$. More formally, the inversely masked subnetwork (i.e., remaining model), $f(x,\tilde{{\bm{m}}}\odot{\bm{\theta}})$, where $\tilde{{\bm{m}}}=1-{\bm{m}}$, should have difficulty expressing $K_{T}$. We define this as the suppression criterion, as it encourages that the remaining model cannot represent knowledge in $K_{T}$. If we find such a disentanglement, we consider that the pretrained model heavily relied on the removed subnetwork to perform a task related to $K_{T}$.
Maintenance   However, only optimizing for suppression leaves the possibility that our method may discover subnetworks that are critical to all expressions of knowledge, or expressions of any coherent sequence of language. As the model should retain most of its initial capacities, we also define maintenance criteria that knowledge-critical subnetworks must follow: (1) they should not affect the model’s original performance on other relational knowledge $K_{C}=K\setminus K_{T}$ (ControlKG), and (2) they should not affect the model’s original language modeling abilities on a standard dataset $D_{LM}$ (ControlLM). We refer to these criteria as maintenance-KG and maintenance-LM respectively.
Sparsity   Finally, we would like the percentage of parameters pruned for the critical subnetwork to be as high as possible to find the parameters that primarily encode the expression of $K_{T}$. There may be irrelevant parameters that are not essential to the expression of $K_{T}$ or $K_{C}$ that do not get pruned from the critical subnetwork if we do not enforce a high sparsity.

### v2

@@ Abstract
Pretrained language models (LMs) encode implicit representations of knowledge in their parameters. However, localizing these representations and disentangling them from each other remains an open problem. In this work, we investigate whether pretrained language models contain various knowledge-critical subnetworks: particular sparse computational subgraphs that can, if removed, precisely suppress specific knowledge the model has memorized. We propose a multi-objective differentiable masking scheme that can be applied to both weights and neurons to discover such subnetworks and show that we can use them to precisely remove specific knowledge from models while minimizing adverse effects on the behavior of the original model. We demonstrate our method on multiple GPT2 variants, uncovering highly sparse subnetworks (98%+ sparsity) that are critical for expressing specific collections of relational knowledge. When these subnetworks are removed, the remaining network maintains most of its initial abilities but struggles to represent the suppressed knowledge. [^1] The code is made available at https://github.com/bayazitdeniz/know-subnet
Discovering Knowledge-Critical Subnetworks
in Pretrained Language Models
Deniz Bayazit, Negar Foroutan, Zeming Chen, Gail Weiss, Antoine Bosselut EPFL {deniz.bayazit,antoine.bosselut}@epfl.ch
@@ 1 Introduction
Figure 1: Knowledge-critical subnetworks are necessary for expressing target knowledge triplets (TargetKG) in LMs. When removed, the remaining model no longer expresses the specific triplets, but maintains its ability to express other relational knowledge (ControlKG) and its language modeling abilities (ControlLM).
Large-scale language models (LLMs) encode large amounts of relational knowledge Petroni et al. (2019); Carlini et al. (2023); Liu et al. (2023), which they transfer to successfully adapt to downstream tasks Wang et al. (2019b, a). Following this success, considerable research focuses on better understanding the extent to which LLMs capture this knowledge Liu et al. (2019a); Safavi and Koutra (2021); Da et al. (2021); Huang et al. (2022). In these works, relational triplets (e.g., (car, IsA, vehicle)) are converted to natural language (e.g., ‘‘A car is a vehicle.’’) before being presented to a model. Key tokens in these input sequences are masked, and the model demonstrates its knowledge of the relations by recovering these tokens.
With the body of work studying LLMs as knowledge bases, a subset of works focuses on where and how this knowledge may be encoded by the models that capture it. The answer to these questions could potentially facilitate the development of more effective finetuning methods, which can be useful for rectifying factual errors made by language models, updating models with evolving knowledge, and preventing ethically undesirable behavior.
Considerable work in model probing Belinkov and Glass (2019); Durrani et al. (2020); Antverg et al. (2022); Belinkov (2022) and mechanistic interpretability Geva et al. (2021, 2022b, 2022a) explores these questions, discovering hidden representations, neurons, and layers that are responsible for the expression of knowledge from these systems. However, these works typically do not localize the knowledge accessing behavior to individual parameters. Another line of work in model editing explores whether knowledge in the model can be changed De Cao et al. (2021); Dai et al. (2022); Hase et al. (2023b); Mitchell et al. (2022a, b); Meng et al. (2022, 2023); Hase et al. (2023a); Gupta et al. (2023); Jang et al. (2023); Chen et al. (2023). However, the goal of these methods is also typically not to precisely localize the parameters responsible for expressing knowledge, but instead to broadly edit model parameters such that a new desired behavior overwrites the model’s preference for the old one.
In this work, we hypothesize that any piece of relational knowledge expressed by a language model is encoded by a limited subset of its parameters. We search for these parameters by identifying sparse subnetworks that, when removed, suppress the model’s ability to express the knowledge of interest while not affecting other abilities of the model. As the model cannot express target knowledge without these subnetworks, we refer to them as knowledge-critical. In Figure 1, we illustrate this concept – when the weights marked with a red cross are removed from the original network, the expression of the triplet (cafe, IsA, restaurant) is suppressed, whereas other triplets are not.
@@
To discover knowledge-critical subnetworks, we propose training differentiable masks over weights or neurons of the original pretrained model, such that the mask can identify and remove a knowledge-critical subnetwork for the targeted knowledge graph. Specifically, we train the mask to: (1) suppress the expression of the target knowledge triplets, (2) maintain the ability to express generic relational knowledge and language, and (3) remove only a minimal subset of weights. After training, the remaining pruned model can no longer express the target knowledge, but maintains its performance on other behaviors, thereby identifying the knowledge-critical subnetwork as the masked portion of the original model.
@@
Our results — across multiple target knowledge graphs (constructed from WordNet and ConceptNet) and LLMs at multiple scales (from the family of GPT2 models) — show that weight masking consistently identifies sparse subnetworks (an average sparsity of $\sim$98.6%) that satisfy our objectives. When these subnetworks are removed, the remaining model’s perplexity on the target knowledge associated with the subnetwork largely increases (an average relative perplexity increase of 253% - 5589% for different GPT2 models), indicating that the expression of the target knowledge is successfully suppressed. However, the remaining network’s ability to model generic relational knowledge and natural language negligibly changes. Finally, in a study on CommonsenseQA, we show that once these subnetworks are removed, models finetuned using parameter-efficient methods struggle with questions that require the knowledge encoded by the removed subnetworks.
@@ 2 Related Work
Our work builds on prior research that demonstrates the knowledge memorization abilities of large language models (LLMs; Carlini et al., 2021; AlKhamissi et al., 2022). Multiple studies have shown that LLMs encode various types of knowledge Liu et al. (2019a); Chen and Gao (2022); Safavi and Koutra (2021); Huang et al. (2022). In these works, parametric knowledge in LLMs is typically expressed by conditioning on a natural language context to complete or infill a sequence that expresses the knowledge Petroni et al. (2019); Jiang et al. (2020); Shin et al. (2020); Cao et al. (2021a); Zhong et al. (2021); Qin and Eisner (2021); Liu et al. (2023); Yu et al. (2023). Other methods also fine-tune models to create an interface to parametric knowledge Bosselut et al. (2019); Roberts et al. (2020); Jiang et al. (2021); Hwang et al. (2021). In contrast, our work investigates where knowledge is encoded by LLMs and localizes the critical subnetworks for expressing these facts.
Function-Specific Subnetworks
@@
Methodologically, our work draws inspiration from studies that identify task-specific subnetworks in neural networks. Perhaps most known, Frankle and Carbin (2019) propose the Lottery Ticket Hypothesis, showing that learned subnetworks could achieve test accuracy similar to that of original networks. Other works prune subnetworks for the purpose of efficient finetuning Mallya et al. (2018); Zhao et al. (2020); Sanh et al. (2020); Guo et al. (2021), or identifying function-specific subnetworks Cao et al. (2021b); Sanh et al. (2020); Zhang et al. (2021); Csordás et al. (2021). Identifying function-specific subnetworks also leads to useful applications, such as disentangling representations to reduce model susceptibility to spurious correlations Zhang et al. (2021), probing models for linguistic properties Cao et al. (2021b); De Cao et al. (2020), identifying and removing a toxic behavior or bias Li et al. (2024); Chintam et al. (2023), and finding subnetworks specialized for different languages Foroutan et al. (2022). Most similar to our work is that of Ren and Zhu (2022), which learns coarse subnetworks that encoded large portions of ConceptNet. We also adopt a differentiable weight masking scheme, but use it to identify highly sparse subnetworks critical for particular expressions of knowledge.
Mechanistic Interpretability
@@
Mechanistic interpretability tackles the problem of understanding model behavior by reverse-engineering computations performed by transformer models. Elhage et al. (2021) discovered algorithmic patterns and frameworks in simplified transformer models. Following this, researchers discovered induction heads Olsson et al. (2022), i.e., specific attention heads involved in in-context learning in LLMs. Similarly, with interventions on attention and MLP sublayers, Geva et al. (2023) identified critical points where the model propagates information, as well as the internal mechanism for attribute extraction. Other work focuses on knowledge tracing and localization in model parameters for the goal of model editing Dai et al. (2022); Meng et al. (2022, 2023); Gupta et al. (2023); Hernandez et al. (2024). Activation patching with corrupted tokens Meng et al. (2022) or corrupted prompts Wang et al. (2023) use causal intervention to identify activations responsible for flipping the model’s output. In contrast, our work focuses on preserving the original model to precisely locate individual model weights responsible for expressing a given set of target knowledge without counterfactuals. Our work is closer to path patching Goldowsky-Dill et al. (2023) and automatic circuit discovery Conmy et al. (2023) to localize behaviors in network subgraphs but focuses specifically on identifying subnetworks associated with knowledge relationships. Our work is also similar to Lo et al. (2024), which shows that models can re-learn removed concepts via neurons. In contrast, we focus on individual parameter pruning.
3 Background & Considerations
@@
To find a knowledge-critical subnetwork in a pretrained language model, we learn a differentiable parameter mask (§4) using a prediction task where the LM is prompted for relational knowledge.
Prompting LMs with KGs
We define a global relational knowledge graph (KG) as the set of knowledge triplets,
$K=\{(h_{1},r_{1},t_{1}),...(h_{n},r_{n},t_{n})\}$
@@
where $h$ and $t$ are head and tail entity nodes, respectively, and $r$ is the relation that holds between the two entities. To input relational knowledge to an LM, triplets are verbalized using a natural language template. For example, the triplet (house, IsA, building), can be verbalized with the template ‘‘{article} {$h$} is {article} {$t$}’’ as ‘‘A house is a building.’’ A typical way to prompt for knowledge is to mask the tail entity ‘‘A house is a ___’’ Petroni et al. (2019). To approximate an autoregressive model’s confidence in a given triplet, we compute a distribution over the missing token and calculate the perplexity of the model for the correct token building.
Differentiable Weight Masking for Function-Specific Parameter Search
@@
To localize parameters that are critical for modeling specific knowledge, we learn a binary mask over each network parameter. For a language model $f(x,{\bm{\theta}}){}$ with pretrained parameters ${\bm{\theta}}$ that takes as input $x$, we learn a set of binary parameters ${\bm{m}}\in\{0,1\}^{|{\bm{\theta}}|}$ that is element-wise multiplied with the frozen ${\bm{\theta}}$, such that our subnetwork is formulated as $f(x,{\bm{m}}\odot{\bm{\theta}}){}$. Similar to other binary mask learning methods Cao et al. (2021b); Sanh et al. (2020), our method models each parameter mask ${\bm{m}}_{i}$ with the concrete (i.e., Gumbel-Softmax) distribution, a differentiable approach to learn continuous mask scores ${\bm{s}}_{i}\in[0,1]$ from real-valued parameters ${\bm{l}}_{i}\in{\mathbb{R}}$ Maddison et al. (2017); Jang et al. (2017):
@@
where ${\mathcal{U}}_{1},{\mathcal{U}}_{2}\sim{\mathcal{U}}(0,1)$ and $\sigma$ is a sigmoid function. We use the approach of Csordás et al. (2021), which uses a straight-through estimator that thresholds the continuous score Bengio et al. (2013):
@@
Mask Granularity
Discovering subnetworks requires selecting the granularity of the parameter mask, reflecting the granularity at which we hypothesize separable knowledge representations can be discovered in the model. Most prior work selects neurons (Elhage et al., 2022) or layers (Zhou et al., 2023) as the basic structural unit for localizing model behaviors. While these representations have been shown to encode knowledge behaviors (Dai et al., 2022; Lo et al., 2024), they are perhaps too broad for reliably disentangling specific knowledge, as they are typically polysemantic (i.e., they jointly encode multiple behaviors; Olah et al., 2020). Conversely, localizing knowledge representations as an unconstrained combination of individual parameters is likely more separable, but may be noisy, as many parameters may be largely redundant, and individual parameters may suffer from overfitting. With no clear choice, in this work, we explore both parameter-level and neuron-level masking to provide complementary insights for mechanistic knowledge localization.
@@ 4 Methodology
This section defines our methodology for discovering knowledge-critical subnetworks using differentiable weight or neuron masking.
Notation
We define a subnetwork as in §3: $f(x,{\bm{m}}\odot{\bm{\theta}})$, where ${\bm{\theta}}$ is the set of parameters of the network $f$ and ${\bm{m}}$ is the mask over a portion of that network’s parameters. To learn a mask over neurons, we jointly mask all the weights connecting to the same neuron. We assume a target set of knowledge $K_{T}\subset K$ (TargetKG) for which we want to identify the critical parameters.
@@ 4.1 Knowledge-Critical Subnetworks
Our goal is to find knowledge-critical subnetworks: the essential parameters to express a given set of target knowledge. When knowledge-critical subnetworks are removed, the expression of the target triplets should be suppressed, and the expression of irrelevant triplets should be unaffected.
Suppression
For $f(x,{\bm{m}}\odot{\bm{\theta}})$ to be critical in expressing $K_{T}$, its removal from the original network should also suppress the model’s ability to express the knowledge in $K_{T}$. More formally, the inversely masked subnetwork (i.e., remaining model), $f(x,\tilde{{\bm{m}}}\odot{\bm{\theta}})$, where $\tilde{{\bm{m}}}=1-{\bm{m}}$, should have difficulty expressing $K_{T}$. We define this as the suppression criterion, as it encourages that the remaining model cannot represent knowledge in $K_{T}$. If we find such a disentanglement, we consider that the pretrained model heavily relies on the removed subnetwork to perform a task related to $K_{T}$.
Maintenance
However, if only optimized for suppression, our method may discover subnetworks that are critical to all expressions of knowledge, or all expressions of coherent sequences of language. As the model should retain its initial capacities, we also define maintenance criteria for knowledge-critical subnetworks. They should: (1) not affect the model’s ability to express other relational knowledge $K_{C}=K\setminus K_{T}$ (ControlKG), and (2) not affect the model’s original language modeling abilities (ControlLM). These criteria are referred to as maintenance-KG and maintenance-LM, respectively.
Sparsity
Finally, we aim to keep the knowledge-critical subnetwork as sparse as possible to discover the parameters that predominantly encode the expression of $K_{T}$. Without imposing a high sparsity level, parameters unrelated to the expression of $K_{T}$ or $K_{C}$ might persist within the subnetwork.
Knowledge Graph # triplets # heads # tails # rels GPT-2 PPL
Small Med Large XL
WordNet ControlKG train 9751 9707 2709 1 63.6 32.8 27.4 24.5
ControlKG val. 50 50 50 1 73.2 37.5 31.3 30.8
building 11 11 11 1 51.9 - - -
communication 16 16 9 1 96.3 65.2 69.2 59.8
change 13 13 13 1 109.7 - - -
statement 16 16 16 1 170.2 - -
location 19 19 7 1 198.0 119.0 125.5 81.4
representation 12 12 12 1 210.7 106.8 108.7 85.0
magnitude 12 12 7 1 299.9 - - -
ConceptNet ControlKG train 5455 2898 2129 16 373.0 - - -
ControlKG val. 606 522 482 16 172.3 - - -
fruit 36 11 37 12 381.6 - - -
sun 36 11 36 12 387.5 - - -
swimming 40 14 40 15 517.8 - - -
Table 1: Statistics on sampled KGs and their verbalization. The graph statistics show the number of triplets and the unique number of heads, tails, and relations. The average perplexity is calculated with the gold tail token cross-entropy loss. The perplexity for certain KGs in the Medium, Large and XL columns are not included as we do not evaluate them in our study on model scale.